# “Talking Past Each Other”: Issue Ownership and Microtargeting in Swiss Online Political Ads

### v1

@@ Abstract
Switzerland’s unique system of direct democracy, characterized by frequent popular referenda, provides a critical context for studying the impact of online political advertising beyond standard electoral cycles. This paper presents a large-scale, data-driven analysis of 40k political ads published on Facebook and Instagram in Switzerland between 2021 and 2025. Despite a voting population of only 5.6 million, the ad campaigns were significant in scale, costing CHF 4.5 million and achieving 560 million impressions. This study shows that political ads are used not only for federal elections, but also to influence referenda, where greater exposure to “pro-Yes” advertising correlates significantly with approval outcomes. The analysis of microtargeting reveals distinct partisan strategies: centrist and right-wing parties predominantly target older men, whereas left-wing parties focus on young women. Furthermore, significant region-specific demographic variations are observed even within the same party, reflecting Switzerland’s strong territorial divisions. Regarding content, a clear pattern of “talking past each other” is identified: in line with issue ownership theory, parties avoid direct debate on shared issues, preferring to promote exclusively owned topics. Finally, it is demonstrated that these strategies are so distinct that an ad’s author can be predicted using a machine learning model trained exclusively on its audience and topic features. This study sheds light on how microtargeting and issue divergence on social platforms may fragment the public sphere and bypass traditional democratic deliberation.
@@ 1 INTRODUCTION
The efficacy of direct democracy is often questioned in academic literature, with some analyses finding it to be ineffective or even counterproductive in practice. Critics often highlight the potential for a “tyranny of the majority”, the influence of wealthy special interest groups, and the risk posed by citizen incompetence regarding complex policy issues [warner1995direct, gerber2011populist]. However, Switzerland is often cited as an exception to this criticism, with its political system considered a model of success [BrunoFrey]. This Swiss model is not a simple and universally applicable template; rather, it has developed within a specific and unique historical and institutional setting. It evolved "bottom-up" from cantonal traditions, not being imposed by a central authority [Linder2010].
Swiss citizens vote on national issues approximately four times per year, participating in far more referendums and initiatives than citizens of any other country [Serdült2014]. This high degree of participation is argued to have numerous positive social and economic effects. These include empowering citizens, making procedures more appealing and imposing fiscal discipline on public spending [BrunoFrey, STUTZERALOISBRUNO]. Further research connects direct democratic rights directly to improved economic performance [FeldLarsSaviozMarcel] and higher levels of subjective well-being, or happiness, among citizens [FreyBrunoStutzerAlois2000].
@@
Even within the Swiss context, the positive portrayal of direct democracy has its detractors. Some scholars have highlighted problematic aspects, including “voting fatigue” caused by the large number of decisions, the complexity of certain ballot measures, and the system’s ability to impede social and political reforms [BLONDELJEAN]. Furthermore, the “double majority” rule, which requires a majority of both citizens and cantons, has been criticized for giving disproportionate power to small, rural cantons. This rule enables controversial initiatives to pass despite lacking broad popular support in major population centers [Papadopoulos2001].
Beyond these long-standing debates, the proliferation of digital media introduces a new and critical challenge to direct democracy. In fact, online political advertising has long been considered controversial and potentially capable of undermining democratic processes. This idea is based on studies highlighting its effectiveness in manipulating voters, fueling cynicism, and polarizing the electorate [JamiesonKathleenHall]. The advent of microtargeting, which uses vast datasets to tailor persuasive messages to narrow, psychologically profiled segments of the electorate, significantly amplifies these threats and introduces new challenges [ZuiderveenBorgesiusFrederik]. In fact, personalized political ads are found to be more effective at persuasion than generic, non-personalized messages [SimchonEdwards, TappinWittenberg].
Active democracy is founded on the idea that citizens are well-informed and capable of debating and deciding on complex policy issues as part of shared public deliberation [BLONDELJEAN]. If microtargeting can effectively sway support with personalized, often emotional, and non-public messages, it threatens to bypass this traditional public sphere. Could new forms of digital propaganda potentially hinder effective democratic participation? What happens when microtargeting and opaque online political ad campaigns are exploited in a country that prides itself on direct democracy as a source of national identity?
These questions require a data-driven approach to assess the normative impact and understand the scale and scope of this phenomenon. This paper aims to present a comprehensive analysis of political advertising on social media in Switzerland, providing an empirical foundation for a critical debate. The main research questions are the following:
  • RQ1 Microtargeting
@@
    b: To what extent can ads influence referendum outcomes?
@@
    b: Is the content of an ad a reliable predictor of its author?
These research questions are addressed by collecting and analyzing 40,000 political ads published on Facebook and Instagram from 2021 to 2025. Through an analysis of demographic microtargeting and textual content, this study highlights the differences and similarities among Switzerland’s major political parties and linguistic regions. This is the first research in the literature to employ statistical models and techniques to examine such a large volume of online political ads published in Switzerland.
2 BACKGROUND AND RELATED WORK
2.1 Swiss Context
In order to understand the deep cultural roots of direct democracy in Swiss institutions and society, it is necessary to start with the very foundations of the Swiss Confederation. The modern federal state was established in 1848 as a direct result of a brief civil war, known as the Sonderbund War. The new state faced the challenges of integrating minorities and managing the country’s deep “cross-cutting cleavages” [Linder2010]. Switzerland was and remains a mosaic:
Linguistic. Since unification, there have been four national languages: German, French, Italian, and Romansh. In 2023, 61.4% of the population is native German speaker; 22.6% are native French speakers; 7.7% are native Italian speakers; and 0.5% are native Romansh speakers [FSOLang].
Religious. According to the Swiss National Survey [FSOrelsurvey], 35.6% of the population was religiously unaffiliated in 2023. The remaining population was divided as follows: 30.7% Catholic, 19.5% Swiss Protestant, 5.8% other Christian denominations, and 6% Muslim.
Geographic. Although rural municipalities cover 57% of the country’s total area, in 2025 they are home to only 35% of the population. In contrast, urban areas account for just 17% of the total area but are home to 65% of the population and 76% of the workforce [FSOGeo].
Economic. Cantons with major urban centers, such as Zurich and Basel-Stadt, have an exceptionally high economic output compared to other cantons. For instance, the nominal GDP per capita in Basel-Stadt was about CHF 210,000 in 2022, compared to the Swiss national average of around CHF 90,000 and the canton of Uri’s average of CHF 58,000 [FSO_GDP].
The federal state’s guarantee of municipal and cantonal autonomy with regard to taxation, education, and infrastructure has played an important role in preserving cohesion [Sager_Zollinger_2011]. Alongside this federal structure, the system of direct democracy has been the most important political compromise made to maintain cohesion in diversity since the unification in 1848. For this reason, Swiss direct democracy has often been interpreted by scholars as a mechanism for minority protection and power-sharing, specifically designed to integrate the country’s diverse segments [Linder2010, Disset2010]. The legal instruments available to citizens and enshrined in the Swiss Federal Constitution include the compulsory referendum, non-compulsory (facultative) referendum, and the popular initiative [Kriesi2005, Papadopoulos2001].
Swiss direct democracy operates within a historically highly fragmented, multipolar party system that reflects the nation’s underlying divisions. In 2025, the Federal Assembly remains highly multipolar, with no single party holding a majority. Following the 2023 federal elections, the national-conservative SVP (Swiss People’s Party) affirmed its position as the largest party with 27.9% of the vote. The center-left SP (Social Democratic Party) followed at 18.3%. The political center is split between the liberal, center-right FDP (FDP.The Liberals) at 14.3% and the Christian democratic Die Mitte (The Center) at 14.1%. The environmentalist bloc, which saw historic gains in 2019, is divided between the left-leaning GRÜNE (Green Party) at 9.8% and the centrist GLP (Green Liberal Party) at 7.6% [FSO_Elections2023]. Finally, it should be mentioned that in Switzerland, in 2025, out of a population of approximately 9.1 million inhabitants, only about 5.6 million are eligible to vote.
2.2 Online political advertising
The influence of social media platforms on political opinion formation and voting behavior remains a topic of debate in the scientific literature. For instance, large-scale experiments have demonstrated that algorithmically delivered social messages can measurably increase voter turnout [bond201261]. It has been proven that curated search rankings can also potentially shift the preferences of undecided voters [Epstein]. Other studies have examined the impact of online communication on politically sensitive topics. For example, Papakyriakopoulos4 investigated the role of demagoguery in civic engagement on Reddit, while LeeYu-Hao examined participation in online activism, or “slacktivism”, in donation campaigns. Memes and other new forms of communication have also been studied as a mean of political propaganda on social media [WilliamsNagel, Mejova, Leon].
This online political context intersects with the underlying business model of social media platforms, which is predominantly based on advertising [ANDERSON201541]. The economic logic incentivizes the continuous algorithmic optimization of content delivery systems to maximize user engagement and, consequently, advertising revenue [O’Reilly_Strauss_Mazzucato_2024, RanDezhi]. The exploitation of these highly optimized, profit-driven systems for the dissemination of political messages is, on the one hand, a source of contemporary public debate and academic research [DommettKatharine, AliSapiezynski]. On the other hand, the effectiveness of recommendation systems has also made online ads much more effective [ChuVliegenthart] and, as a result, led to a significant increase in their use [Sosnovik]. In this context, Meta’s platforms (Facebook and Instagram) stand out as among the most effective in influencing public opinion [KarpfDavid]. madrigal2017facebook argues that although Facebook’s potential impact on elections had been evident for years, it reached its peak during the 2016 presidential election, with \$1.4 billion spent on online ads during the campaign. However, hundreds of millions of dollars of dark ads paid for by foreign governments to influence elections should be added to this number [madrigal2017facebook]. In this regard, it has been proven that Russia attempted to manipulate the 2016 U.S. presidential election by running political ads on Facebook [entous2017russian, EmilioFerrara]. MatiasNathan designed a software-supported approach for auditing, which uses coordinated volunteers to analyze political advertising policies enacted by Facebook and Google during the 2018 midterm U.S. election. Other studies on Facebook political ads include anti-vaccination ads during COVID-19 [MejovaKalimeri], populist advertising for the European Parliamentary election in 2019 [CapozziThin], Italian political propaganda on migrants issue [capozzi2021clandestino, 10.1007/978-3-030-60975-7_4], or far-right misinformation spread by the populist Spanish party VOX [CanoLorena]. Many studies examined the spread and impact of political ads in Switzerland, with the earliest focusing on traditional newspapers [GerberBühlmann, ZumofenGuillaume]. Other studies have focused on referendum campaigns shifting from newspapers to social media [FischerMichaela], or politicians’ use of social media [ReveilhacMorselli].
3 DATA
This study relies on data from two main sources: the Swiss Federal Statistical Office [^1] https://abstimmungen.admin.ch/en/overview and the Meta Ad Library API [^2] https://www.facebook.com/ads/library/api/. The Federal Statistical Office provides detailed statistics related to referenda, including results by canton, participation percentages, and full descriptions of referendum questions. All data collected and annotated [^3] Zenodo anonymized URL and the code written [^4] GitHub anonymized URL are publicly available.
3.1 Meta Ad Library
The Meta Ad Library API discloses advertisements about social issues, election, or politics that were delivered on Facebook or Instagram during the past seven years. Although the API allows for specifying search parameters, such as limiting results to a specific country, all queries must contain keywords or author names. The following information can be extracted for each advertisement: the publication date; the end time of the campaign; the textual content; and the number of impressions (provided as an interval). The term “impressions” refers to non-unique visualizations. It is also possible to access the campaign expenditure (again presented as an interval) and the audience’s demographic profile, specifically their gender (male, female, or unknown) and age (categorized into seven buckets). In line with existing literature [capozzi2021clandestino], the midpoint of the interval’s boundaries is computed for values presented as ranges (expenditure and impressions). For unbounded ranges, only the defined boundary is used.
The collection is initiated using a broad set of election-related keywords (see Appendix Table A1) in German, French, and Italian due to the API’s requirement for keyword- or author-based searches. This process yielded an initial corpus of 40,044 ads (22,494 from German keywords, 10,794 from French, and 6,756 from Italian) published between January 1, 2021, and September 30, 2025. The API’s country filter includes advertisements displayed in Switzerland, even if they are intended for a different audience. To create a Swiss-centric dataset, ads are further filtered to include only those for which more than 50% of impressions come from users based in Switzerland. This reduces the dataset to 34,559 ads from 2,246 unique pages, which received an estimated 560 million impressions and cost over CHF 4.5 million.
Although the API provides a page name (username) for each ad, this identifier is frequently ambiguous or does not directly correspond to a particular political party, even when the content of the page clearly promotes a party’s agenda. As shown in Figure 1, the 2023 federal elections represent the most significant activity peak in the dataset. For this reason, the 1,224 pages that published at least one ad in the 30 days preceding the federal election are manually labeled to establish whether they are related to a specific party. This process identified 943 pages that could be directly attributed to a political party, while the remaining 281 pages belong to entities such as media organizations (newspapers and daily newspapers), NGOs, neighborhood committees, and non-partisan referendum groups. Of the 943 political pages, 17 major parties and smaller entities (e.g., municipal parties and civic lists) are identified. However, only six parties — SVP, SP, FDP, Die Mitte, GRÜNE, and GLP — play a dominant role in the advertising landscape, accounting for 95% of all impressions. Collectively, these parties also secured 91.84% of the votes in the 2023 federal elections. Tables  A2,  A3, and  A4 provide detailed statistics on impressions, volume, and ad expenditure broken down by party and language.
3.2 Content Annotation
3.2.1 Referenda
3.2.2 Federal Election
The second analysis focuses on the 2023 federal elections. Unlike a referendum, which addresses one or a few political issues, federal elections involve political parties promoting broader agendas. Therefore, it is necessary to not only identify election-relevant ads but also determine the specific topic(s) they cover. Consistent with the previous task, only ads with at least 10 words published within 30 days before the federal election are selected. The model checkpoint and the API query date are the same of the previous task. The model’s output includes a Boolean parameter identifying whether the ad is relevant to the federal election, as well as up to three English keywords related to the ad’s topic. This initial extraction process yielded a large set of 739 unique keywords, which require consolidation into fewer categories. To this end, each keyword is encoded into a semantic vector using OpenAI’s state-of-the-art text-embedding-3-large model. These vectors are then clustered into 25 topics using the K-Means clustering algorithm [mcqueen1967some]. To ensure reliability, qualitative validation is performed on a random subset of 100 ads. In this review, human annotation agreed with model-generated keywords in 96% of the cases.
4 RESULTS

### v2

# The direct democracy paradox: Microtargeting and issue ownership in Swiss online political ads

@@ Abstract
Political advertising on social media has fundamentally reshaped democratic deliberation, playing a central role in electoral campaigns and propaganda. However, its systemic impact remains largely theoretical or unexplored, raising critical concerns about institutional fairness and algorithmic transparency. This paper provides the first data-driven analysis of the relationship between direct democracy and political advertising on social media, leveraging a novel dataset of 40,000 political ads published on Meta in Switzerland between 2021 and 2025. Switzerland’s system of direct democracy, characterized by frequent referenda, provides an ideal context for examining this relationship beyond standard electoral cycles. The results reveal the sheer scale of digital campaigning, with 560 million impressions targeting 5.6 million voters, and suggest that greater exposure to “pro-Yes” advertising significantly correlates with referendum approval outcomes. Demographic microtargeting analysis suggests partisan strategies: Centrist and right-wing parties predominantly target older men, whereas left-wing parties focus on young women. Regarding textual content, a clear pattern of “talking past each other” is identified; in line with the issue ownership theory, parties avoid debating shared issues, preferring to promote exclusively owned topics. Furthermore, the parties’ strategies are so distinctive that a machine learning model trained only on audience and topic features can accurately predict the author of an advertisement. This article highlights how demographic microtargeting, issue divergence, and tailored messages could undermine democratic deliberation, exposing a paradox: Referenda are designed to be the ultimate expression of the popular will, yet they are highly susceptible to invisible algorithmic persuasion.
Keywords political advertising, microtargeting, elections, Facebook, Instagram
@@ 1 Introduction
Online political advertising has long been considered controversial and potentially capable of undermining democratic processes, primarily due to its effectiveness in manipulating voters, fueling cynicism, and polarizing the electorate [36]. This reality stands in stark contrast to the principles of active democracy, which is founded on the idea that citizens are well-informed and capable of debating and deciding on complex policy issues as part of shared public deliberation [9]. The threat to this public sphere is significantly amplified by the advent of microtargeting, which uses vast datasets to tailor persuasive messages to narrow, psychologically profiled segments of the electorate [63]. In fact, multiple studies have found that such personalized political ads are more effective at persuading than generic, non-personalized messages [55, 60], while new generative AI technologies promise to render these microtargeting campaigns even more effective [55]. Although the impact of these technologies on candidate elections is well documented [17, 6], their systemic impact on policy-making and direct legislation remains largely unexplored. Does algorithmic microtargeting simply encourage voter participation, or does it fundamentally fragment the debate on key issues? This gap is critical because the risks are amplified in direct democratic systems. Here, a dangerous paradox emerges: citizens operate under the impression that they are making sovereign decisions, yet their preferences may actually be steered by opaque, tailored messaging. If the electorate feels empowered to decide but is unknowingly steered by algorithms, is the “will of the people” still authentic?
To answer this question, Switzerland offers a unique laboratory. Firstly, direct democracy is not just a feature, but a central pillar of the nation’s identity. Swiss citizens vote on national issues approximately four times per year, participating in far more referendums and initiatives than citizens of any other country [Serdült2014]. Secondly, Switzerland is a robust setting for testing the generalizability of cross-cultural findings. Switzerland is a multicultural nation of 9 million people with four official languages, where 41% of permanent residents have a migration background as of 2024.
This article presents a longitudinal, data-driven analysis of a novel dataset of 40,000 political ads published on Facebook and Instagram in Switzerland from 2021 to 2025. Through a combined study of demographic microtargeting and textual content, this research assesses the scale, scope, and normative impact of opaque online political advertising, while isolating differences and similarities among major political parties and linguistic regions. Beyond the specific case study, the article describes a reproducible research pipeline that can be applied to other contexts and democracies, and provides researchers and policymakers with a toolbox to make the digital public sphere more transparent.
2 Background
2.1 Online political advertising
The real impact of social media platforms on political opinion formation and voting behavior remains a topic of debate in the scientific literature. For instance, large-scale experiments have demonstrated that algorithmically delivered social messages can measurably increase voter turnout [10]. Other experiments have shown that curated search rankings can also potentially shift the preferences of undecided voters [23]. The impact of online communication on politically sensitive topics has also been examined. For example, Papakyriakopoulos et al. [50] investigated the role of demagoguery in civic engagement on Reddit, while Lee and Hsieh [39] examined participation in online activism, or “slacktivism”, in donation campaigns. Memes and other new forms of communication have also been studied as a mean of political propaganda on social media [47, 45, 32].
This online political context intersects with the underlying business model of social media platforms, which is predominantly based on advertising [3]. The economic logic incentivizes the continuous algorithmic optimization of content delivery systems to maximize user engagement and, consequently, advertising revenue [O’Reilly_Strauss_Mazzucato_2024, 52]. The exploitation of these highly optimized, profit-driven systems for the dissemination of political messages is, on the one hand, a source of contemporary public debate and academic research [20, 1]. On the other hand, the effectiveness of recommendation systems has also made online ads much more effective [18] and, as a result, led to a significant increase in their use [56].
In this context, Meta’s platforms, Facebook and Instagram, are among the most effective at influencing public opinion [37] and accurately identifying users’ demographic information [35]. Madrigal [41] argues that although Facebook’s potential impact on elections had been evident for years, it reached its peak during the 2016 presidential election, with \$1.4 billion spent on online ads during the campaign. However, hundreds of millions of dollars of dark ads paid for by foreign governments to influence elections should be added to this number [41]. In this regard, it has been proven that Russia attempted to manipulate the 2016 U.S. presidential election by running political ads on Facebook [22, 21]. Matias et al. [43] designed a software-supported approach for auditing, which uses coordinated volunteers to analyze political advertising policies enacted by Facebook and Google during the 2018 midterm U.S. election.
During the 2021 German federal elections, a positive correlation was identified between social media advertising and successful election outcomes for candidates [6]. In addition, Bär et al. [5] found that Meta’s algorithm prioritized advertisements from populist parties during the same election campaign. Other studies on Facebook political ads include anti-vaccination ads during COVID-19 [46], populist advertising for the European Parliamentary election in 2019 [17], Italian political propaganda on the migration issue [16, 15], far-right misinformation spread by the populist Spanish party VOX [14]. Many studies examined the spread and impact of political ads in Switzerland, with the earliest focusing on traditional newspapers [GerberBühlmann, 64]. Other studies have focused on referendum campaigns shifting from newspapers to social media [26], or politicians’ use of social media [53].
2.2 Switzerland and direct democracy
@@
Even before the advent of social media, the efficacy of direct democracy was questioned in academic literature, with some analyzes finding it to be ineffective or even counterproductive in practice. Critics often highlight the potential for a “tyranny of the majority”, the influence of wealthy special interest groups, and the risk posed by citizen incompetence regarding complex policy issues [61, 34]. However, Switzerland is often cited as an exception to this criticism, with its political system considered a model of success [28]. The Swiss model is not a simple and universally applicable template; rather, it has developed within a specific and unique historical and institutional setting. It evolved “bottom-up” from cantonal traditions, not being imposed by a central authority [40]. This high degree of direct democracy and participation is argued to have numerous positive social and economic effects, including empowering citizens, making procedures more appealing, and imposing fiscal discipline on public spending [28, 58]. Further research connects direct democratic rights directly to improved economic performance [25] and higher levels of subjective well-being, or happiness, among citizens [27].
@@
Even within the Swiss context, the positive portrayal of direct democracy has its detractors. Some scholars have highlighted problematic aspects, including “voting fatigue” caused by the large number of decisions, the complexity of certain ballot measures, and the system’s ability to impede social and political reforms [9]. Furthermore, the “double majority” rule, which requires a majority of both citizens and cantons, has been criticized for giving disproportionate power to small, rural cantons. This rule enables controversial initiatives to pass despite lacking broad popular support in major population centers [49].
In order to understand the deep cultural roots of direct democracy in Swiss institutions and society, it is necessary to start with the very foundations of the Swiss Confederation. The modern federal state was established in 1848 as a direct result of a brief civil war, known as the Sonderbund War. The new state faced the challenges of integrating minorities and managing the country’s deep “cross-cutting cleavages” [40]. Switzerland was and remains a mosaic:
Linguistic. Since unification, there have been four national languages: German, French, Italian, and Romansh. In 2023, 61.4% of the population is native German speaker; 22.6% are native French speakers; 7.7% are native Italian speakers; and 0.5% are native Romansh speakers [29].
Religious. According to the Swiss National Survey [31], 35.6% of the population was religiously unaffiliated in 2023. The remaining population was divided as follows: 30.7% Catholic, 19.5% Swiss Protestant, 5.8% other Christian denominations, and 6% Muslim.
Geographic. Although rural municipalities cover 57% of the country’s total area, in 2025 they are home to only 35% of the population. In contrast, urban areas account for just 17% of the total area but are home to 65% of the population and 76% of the workforce [30].
Economic. Cantons with major urban centers, such as Zurich and Basel-Stadt, have an exceptionally high economic output compared to other cantons. For instance, the nominal GDP per capita in Basel-Stadt was about CHF 210,000 in 2022, compared to the Swiss national average of around CHF 90,000 and the canton of Uri’s average of CHF 58,000 [8].
The federal state’s guarantee of municipal and cantonal autonomy with regard to taxation, education, and infrastructure has played an important role in preserving cohesion [54]. Alongside this federal structure, the system of direct democracy has been the most important political compromise made to maintain cohesion in diversity since the unification in 1848. For this reason, Swiss direct democracy has often been interpreted by scholars as a mechanism for minority protection and power-sharing, specifically designed to integrate the country’s diverse segments [40, 19]. The legal instruments available to citizens and enshrined in the Swiss Federal Constitution include the compulsory referendum, non-compulsory (facultative) referendum, and the popular initiative [38, 49].
Swiss direct democracy operates within a historically highly fragmented, multipolar party system that reflects the nation’s underlying divisions. In 2025, the Federal Assembly remains highly multipolar, with no single party holding a majority. Following the 2023 federal elections, the national-conservative SVP (Swiss People’s Party) affirmed its position as the largest party with 27.9% of the vote. The center-left SP (Social Democratic Party) followed at 18.3%. The political center is split between the liberal, center-right FDP (FDP.The Liberals) at 14.3% and the Christian democratic Die Mitte (The Center) at 14.1%. The environmentalist bloc, which saw historic gains in 2019, is divided between the left-leaning GRÜNE (Green Party) at 9.8% and the centrist GLP (Green Liberal Party) at 7.6% [59]. Finally, it should be mentioned that, out of a population of approximately 9.1 million in 2025, only about 5.6 million people are eligible to vote.
3 Results
The findings presented in this section are divided into three subsections: demographic microtargeting (RQ1), impact on direct democratic votes (RQ2), and strategic ownership of issues during federal elections (RQ3). The research questions are the following:
  • RQ1 Demographic microtargeting
@@ 3 Results
    • b: To what extent can ads influence the outcome of a referendum?
@@
    • b: Are the content and the microtargeting of an ad reliable predictors of its author?
3.1 Microtargeting
Figure 1 compares each party’s audience demographics to the “party complementary” profile, which consists of all advertisements not associated with that party. Since RQ1.a does not concern regional differences, ads from all languages are combined for this analysis.
Overall, centrist and right-wing parties are more oriented towards a male audience, particularly the SVP ($36\%$). On the other hand, the GRÜNE and the SP are more female-oriented, with the GRÜNE at $27\%$ and the SP at $13\%$. In terms of age reach, the SVP and, to a lesser extent, Die Mitte tend to attract an older audience, whereas the GRÜNE and the GLP primarily attract people in the 25–34 age group. Furthermore, the GRÜNE is the only party that extensively reaches the 18–24 age group, yet it underrepresents the over-45 age group more than any other party.
Figure 1: Comparison of demographic reach by age and gender. Each party’s audience is compared against the distribution of all other ads (Party Complementary). Gender bias is measured using the male-to-female Odds Ratio (OR), and age distribution shifts are measured using the Wasserstein Distance (WD).
Figure 2: Language-specific age distributions by party. Distribution of audience age for each party compared to all other ads (Party Complementary) across German, French, and Italian cantons. Shaded areas represent the standard deviation derived from bootstrapping. Differences in distributions are reported as the Wasserstein Distance (WD).
Figure 2 shows the age distribution of the audience reached by the parties, compared to the party complementaries, broken down by language group. The shaded area represents the standard deviation of the audience reached for each age group. The distribution is build using a bootstrapping procedure similar to that described for Figure 3. Significant differences between parties within the same canton can be observed, as well as regional variations within parties. While the SVP focuses on an audience over 45, especially in Ticino, Die Mitte focuses on an older audience in French-speaking cantons only. The SP, in German-speaking cantons, has a WD of almost 0 (no differences with the complementary party’s audience), while in French-speaking cantons, it targets a much more over-45 audience. In Ticino, however, it does the opposite, underrepresenting the over-35 audience and focusing on the 18-24 age range. The GLP, on the other hand, focuses heavily on the 25-34 age range in French-speaking cantons, while in Ticino, the reach is broader (25-45). With the exception of the GRÜNE and the SVP, all parties in the German-speaking cantons show fewer differences compared to their counterparts. In summary, parties differ in their demographic reach (RQ1.a), with evident internal variations emerging across different linguistic regions (RQ1.b).
Figure 3: Male-to-female targeting bias across linguistic regions. Odds ratios comparing the gender distribution of each party’s ads against the complementary set. Values are bootstrapped means, with 1.0 representing gender balance, while a value greater than 1.0 indicates a shift toward a more male audience. Significant results, where the 95% confidence interval does not include the baseline of 1.0, are indicated by an asterisk (*).
3.2 Referenda